# The future of cosmological likelihood-based inference: accelerated high-dimensional parameter estimation and model comparison

### v1

@@ Abstract
We advocate for a new paradigm of cosmological likelihood-based inference, leveraging recent developments in machine learning and its underlying technology, to accelerate Bayesian inference in high-dimensional settings. Specifically, we combine (i) emulation, where a machine learning model is trained to mimic cosmological observables, e.g. CosmoPower-JAX; (ii) differentiable and probabilistic programming, e.g. JAX and NumPyro, respectively; (iii) scalable Markov chain Monte Carlo (MCMC) sampling techniques that exploit gradients, e.g. Hamiltonian Monte Carlo; and (iv) decoupled and scalable Bayesian model selection techniques that compute the Bayesian evidence purely from posterior samples, e.g. the learned harmonic mean implemented in harmonic. This paradigm allows us to carry out a complete Bayesian analysis, including both parameter estimation and model selection, in a fraction of the time of traditional approaches. First, we demonstrate the application of this paradigm on a simulated cosmic shear analysis for a Stage IV survey in 37- and 39-dimensional parameter spaces, comparing $\Lambda$CDM and a dynamical dark energy model ($w_{0}w_{a}$). We recover posterior contours and evidence estimates that are in excellent agreement with those computed by the traditional nested sampling approach while reducing the computational cost from 8 months on 48 CPU cores to 2 days on 12 GPUs. Second, we consider a joint analysis between three simulated next-generation surveys, each performing a 3x2pt analysis, resulting in 157- and 159-dimensional parameter spaces. Standard nested sampling techniques are simply not feasible in this high-dimensional setting, requiring a projected 12 years of compute time on 48 CPU cores; on the other hand, the proposed approach only requires 8 days of compute time on 24 GPUs. All packages used in our analyses are publicly available.
@@ 1 Introduction
Differentiability also unlocks gradient-based sampling algorithms, such as Hamiltonian Monte Carlo (HMC, Duane et al., 1987; Neal, 1996) and its variants, which represent a promising avenue to address the challenges of high-dimensional inference. These techniques allow one to efficiently explore the complex parameter spaces that characterize cosmological data analyses. Yet, they crucially rely on having access to the derivatives of the likelihood function with respect to input parameters, which are in general expensive and inaccurate, especially when computed by finite differences. This obstacle is being overcome through the development of fully-differentiable frameworks for cosmological data analysis, which also benefit from the acceleration provided by the same dedicated hardware of ML models.
@@ 2.1 Parameter estimation
Estimating the posterior distribution through its direct evaluation on a grid of parameters is computationally infeasible due to the curse of dimensionality, which becomes a prohibitive factor even in spaces of moderate dimension. Additionally, the likelihood function often involves solving intricate physical models or running detailed simulations, and the posterior landscape may be multi-modal or have complex correlations between parameters. A grid-based evaluation might miss important regions of the parameter space or require an impractically fine grid to capture these features. Sampling algorithms such as MCMC methods, on the other hand, provide a stochastic exploration that can adapt to the shape of the distribution, and allocate computational resources more efficiently by focusing on the regions of parameter space with high posterior density.
Commonly used MCMC algorithms in cosmology include the Metropolis-Hastings (Lewis & Bridle, 2002; Lewis, 2013), affine-invariant ensemble (Goodman & Weare, 2010; Foreman-Mackey et al., 2013), and nested sampling algorithms (Skilling, 2006; Feroz & Hobson, 2008; Feroz et al., 2009, 2019; Handley et al., 2015a, b), which have found widespread use thanks to their clear formulations and publicly available implementations. Another sampling technique is HMC (Duane et al., 1987; Neal, 1996), which is a more efficient algorithm exploiting Hamiltonian dynamics. By leveraging the information coming from the gradients of the likelihood with respect to the input parameters, HMC concentrates sampling in regions of high posterior mass (the so-called typical set), resulting in a more efficient sampler, particularly in high-dimensional settings. A significantly improved version of HMC is NUTS (Hoffman & Gelman, 2014), which overcomes the required tuning of hyperparameters for the numerical integration of Hamilton equations by preventing the sampler from taking U-turns. By exploring the parameter space in a more efficient way, NUTS results in a significantly higher acceptance rate, and has thus found widespread use.
@@ 2.2 Model comparison
Recently, methods to compute the evidence that are agnostic to the sampling technique and only require posterior samples have gained popularity (McEwen et al., 2021; Srinivasan et al., 2024; Rinaldi et al., 2024). Among them, the learned harmonic mean estimator (McEwen et al., 2021) has been shown to provide robust estimates of the Bayesian evidence in a variety of scenarios, both in likelihood- and simulation-based inference (Polanska et al., 2023, 2024; Spurio Mancini et al., 2023). In particular, Polanska et al. (2024) integrated normalizing flows into the learned harmonic mean framework to learn the internal importance sampling target distribution, enhancing the robustness and scalability of the estimator. In contrast to the learned harmonic mean, alternative approaches (Srinivasan et al., 2024; Rinaldi et al., 2024) compute the evidence for a surrogate of the posterior, limiting their accuracy (see Polanska et al. 2024 for further discussion).
@@ 3.2 Differentiable and probabilistic programming:
Modern ML ecosystems, such as JAX and PyTorch, are constructed with this in mind, requiring that each primitive operation within their libraries is coupled with associated gradient rules. In this way, using the chain rule one can construct more complex operations, which are automatically differentiable (Bartholomew-Biggs et al., 2000; Margossian, 2019). By automatic differentiation gradients can then be computed efficiently and accurately for complex operations, replacing numerical approximations of the derivatives, which are typically slow and unstable. Primarily due to its relative maturity, PyTorch provides a larger bank of primitive operations. However, JAX provides a more flexible ecosystem that is functional in nature (Wadler, 1992) and thus well suited to the implementation of general physical models. In JAX complex operations may more easily be expressed, making it the natural ecosystem for the integration of differentiable programming in, e.g. cosmological models (Campagne et al., 2023), cosmological simulations (Li et al., 2024) or underlying mathematical methods, such as spherical harmonic transforms (Price & McEwen, 2024).
Recently, probabilistic programming languages (PPLs) have been constructed leveraging these differentiable programming ecosystems. Through PPLs, statistical operations such as sampling and conditioning can be expressed using a familiar and intuitive Python syntax, whilst also inheriting the differentiability of the underlying ecosystems upon which they are constructed. For instance, PPLs allows one to easily formulate Bayesian hierarchical models, which are widely used in cosmology (see e.g. Mandel et al., 2011; Shariff et al., 2016; Alsing et al., 2016, 2017; Hinton et al., 2019; Porqueres et al., 2021; Mandel et al., 2022; Porqueres et al., 2023; Loureiro et al., 2023; Sellentin et al., 2023).
In this work we use NumPyro (Phan et al., 2019) [^11] https://num.pyro.ai/en/stable/, the NumPy backend for Pyro (Bingham et al., 2019). NumPyro is well-documented and easily interfaces with JAX, thus providing a mature framework to build a gradient-based inference pipeline. We implement the likelihood for the problems considered (see Sect. 4 and Sect. 5) in the NumPyro PPL, which requires a separate likelihood implementation. Critically, the likelihood must be differentiable: CosmoPower-JAX provides emulation of the matter power spectrum that is differentiable, and we couple it with jax-cosmo (Campagne et al., 2023) [^12] https://github.com/DifferentiableUniverseInitiative/jax_cosmo for differentiable cosmological models to simulate observable spectra.
@@ 3.3 Scalable MCMC sampling: NUTS
HMC is a gradient-based sampling algorithm that is scalable to high-dimensional parameter spaces and does not suffer from the low acceptance rate of the standard Metropolis-Hastings algorithm. NUTS further improves on the acceptance rate of HMC by preventing the sampling trajectory from returning to previously-visited regions of the parameter space. Though it requires a higher number of model evaluations, NUTS typically leads to more reliable convergence, and is thus ideal for sampling in the context of Stage IV surveys.
Traditionally, the gradients required for HMC sampling are computed by finite differences, which is computationally costly and inaccurate. Since we have a differentiable pipeline, gradients can be computed by automatic differentiation, providing acceleration and improved accuracy. In the future, new samplers that further improve the convergence rate leading to even more robust inference can be straightforwardly incorporated into this paradigm.
@@ 3.4 Decoupled and scalable Bayesian model comparison: learned harmonic mean
Figure 1: Corner plot for the $37$-dimensional $\Lambda$CDM model, showing the posterior contours obtained with CosmoPower-JAX in red and the concentrated flow with temperature $T=0.8$ in blue.
@@
4 Cosmic shear analysis with 37 and 39 parameters
@@ 4.3 Computational approaches
Second, we compare this “traditional“ approach with our “future” paradigm, which replaces CAMB with the CosmoPower-JAX emulator, rewrites the likelihood in the auto-differentiable language JAX, making use of jax-cosmo, samples the posterior distribution with NUTS implemented in NumPyro, accelerated by automatic differentiation, and computes a robust estimate of the evidence from posterior samples using the learned harmonic mean implemented in harmonic. We run the NUTS chains in parallel on 12 A100 80GB GPUs, collecting 60 chains of 2000 samples each (after 400 samples of warm-up) for both $\Lambda$CDM and $w_{0}w_{a}$. For harmonic, we train a rational-quadratic spline flow (Durkan et al., 2019), consisting of 4 layers and 64 spline bins, on a single GPU. We separate the NUTS samples into a training set the flow is trained on and an inference set that is used for estimation. We use $30\%$ of chains for training and the remaining $70\%$ for inference to compute the evidence on 48 CPUs with temperature $T=0.8$.
@@ 4.4 Results
The results are reported in Table 1(a). The values of the evidence when using CAMB $+$ nested sampling and CosmoPower-JAX $+$ NUTS $+$ harmonic use different implementations of the likelihood (the latter implemented in a probabilistic programming framework), including different normalizations, so their values cannot be compared directly. Nevertheless, the $\log$ BFs are in good agreement between the different approaches, with $0.78\pm 0.79$ for nested sampling and $1.53\pm 0.07$ for harmonic. Note that the evidence values correctly favour the ground truth model considered for the simulated data. While the evidence computed by the naïve flow estimator is also in agreement with a value of $1\pm 6$, notice that its error is considerably larger than the other estimators, as anticipated. While the evidence values computed by the “traditional” and “future” paradigms are in close agreement, computational times are dramatically different, with the “future” paradigm being two orders of magnitude faster, taking 2 days (with only 12 additional minutes required to compute the evidence) instead of roughly 8 months.
The contours for the posterior samples (red) alongside the concentrated flow at $T=0.8$ (blue) used for inference for this analysis are shown in Fig. 1 and Fig. 2 for the $\Lambda$CDM and $w_{0}w_{a}$ models respectively. Note that the flow is contained within the posterior.
@@ 5.2 Results
All numerical results for this higher dimensional analysis are reported in Table 1(b). The high number of parameters and the need to consider multiple probes pose a significant challenge for “traditional” methods, despite being quite realistic for next-generation surveys. We estimated that the computation of the Bayes factor in such a scenario would require about 12 years on 48 CPU cores, making it effectively impossible to perform model selection in a reasonable time.
On the other hand, the proposed “future” paradigm provides an estimate of the Bayes factor based on 40 (45) converged chains for the $\Lambda$CDM ($w_{0}w_{a}$) model, totalling 8 days on 24 GPUs, with only 17 additional minutes required to compute the evidence. In this case, the flow model we consider is constructed from 2 layers and 128 spline bins. The $\log$ Bayes factor computed by harmonic is $1.9^{+0.7}_{-0.5}$, which correctly favours the ground truth model considered for the simulated data. The evidence computed by the naïve flow estimator is $2\pm 73$: its error is very large, as anticipated, rendering this approach unusable in practice.
The contours plots for the posterior samples (red) alongside the concentrated flow at $T=0.8$ (blue) used for inference for this analysis are shown in Fig. 3 and Fig. 4 for the $\Lambda$CDM and $w_{0}w_{a}$ models respectively. Fig. 5 and Fig. 6 show the one-dimensional marginal plots of the posterior samples (red) alongside the concentrated flow at $T=0.8$ (blue), for all the model parameters for $\Lambda$CDM and $w_{0}w_{a}$ respectively. It can be seen that the flow is indeed contained within the posterior as required by the learned harmonic mean.

### v2

@@ Abstract
We advocate for a new paradigm of cosmological likelihood-based inference, leveraging recent developments in machine learning and its underlying technology, to accelerate Bayesian inference in high-dimensional settings. Specifically, we combine (i) emulation, where a machine learning model is trained to mimic cosmological observables, e.g. CosmoPower-JAX; (ii) differentiable and probabilistic programming, e.g. JAX and NumPyro, respectively; (iii) scalable Markov chain Monte Carlo (MCMC) sampling techniques that exploit gradients, e.g. Hamiltonian Monte Carlo; and (iv) decoupled and scalable Bayesian model selection techniques that compute the Bayesian evidence purely from posterior samples, e.g. the learned harmonic mean implemented in harmonic. This paradigm allows us to carry out a complete Bayesian analysis, including both parameter estimation and model selection, in a fraction of the time of traditional approaches. First, we demonstrate the application of this paradigm on a simulated cosmic shear analysis for a Stage IV survey in 37- and 39-dimensional parameter spaces, comparing $\Lambda$CDM and a dynamical dark energy model ($w_{0}w_{a}$). We recover posterior contours and evidence estimates that are in excellent agreement with those computed by the traditional nested sampling approach while reducing the computational cost from 8 months on 48 CPU cores to 2 days on 12 GPUs. Second, we consider a joint analysis between three simulated next-generation surveys, each performing a 3x2pt analysis, resulting in 157- and 159-dimensional parameter spaces. Standard nested sampling techniques are simply unlikely to be feasible in this high-dimensional setting, requiring a projected 12 years of compute time on 48 CPU cores; on the other hand, the proposed approach only requires 8 days of compute time on 24 GPUs. All packages used in our analyses are publicly available.
@@ 1 Introduction
Differentiability also unlocks gradient-based sampling algorithms, such as Hamiltonian Monte Carlo (HMC, Duane et al., 1987; Neal, 1996) and its variants, which represent a promising avenue to address the challenges of high-dimensional inference. These techniques allow one to efficiently explore the complex parameter spaces that characterize cosmological data analyses. Yet, they crucially rely on having access to the derivatives of the likelihood function with respect to input parameters, which are in general expensive and inaccurate, especially when computed by finite differences. This obstacle is being overcome through the development of fully-differentiable frameworks for cosmological data analysis (see e.g. Nygaard et al., 2023; Ruiz-Zapatero et al., 2024; Balkenhol et al., 2024), which also benefit from the acceleration provided by the same dedicated hardware of ML models.
@@ 2.1 Parameter estimation
Estimating the posterior distribution through its direct evaluation on a grid of parameters is computationally infeasible due to the curse of dimensionality, which becomes a prohibitive factor even in spaces of moderate dimension. Additionally, the likelihood function often involves solving intricate physical models or running detailed simulations, and the posterior landscape may be multimodal or have complex correlations between parameters. A grid-based evaluation might miss important regions of the parameter space or require an impractically fine grid to capture these features. Sampling algorithms such as MCMC methods, on the other hand, provide a stochastic exploration that can adapt to the shape of the distribution, and allocate computational resources more efficiently by focusing on the regions of parameter space with high posterior density.
Commonly used MCMC algorithms in cosmology include the fast-slow random-walk Metropolis–Hastings (Lewis & Bridle, 2002; Lewis, 2013), affine-invariant ensemble (Goodman & Weare, 2010; Foreman-Mackey et al., 2013), and nested sampling algorithms (Skilling, 2006; Feroz & Hobson, 2008; Feroz et al., 2009, 2019; Handley et al., 2015a, b; Buchner, 2021), which have found widespread use thanks to their clear formulations and publicly available implementations. Another sampling technique is HMC (Duane et al., 1987; Neal, 1996), which is a more efficient algorithm exploiting Hamiltonian dynamics. By leveraging the information coming from the gradients of the likelihood with respect to the input parameters, HMC concentrates sampling in regions of high posterior mass (the so-called typical set), resulting in a more efficient sampler, particularly in high-dimensional settings. A significantly improved version of HMC is NUTS (Hoffman & Gelman, 2014), which overcomes the required tuning of hyperparameters for the numerical integration of Hamilton equations by preventing the sampler from taking U-turns. NUTS results in a more efficient exploration of parameter space and faster convergence, and has thus found widespread use.
@@ 2.2 Model comparison
Recently, methods to compute the evidence that are agnostic to the sampling technique and only require posterior samples have gained popularity (Dickey, 1971; Trotta, 2007; Heavens et al., 2017; Jia & Seljak, 2020; McEwen et al., 2021; Srinivasan et al., 2024; Rinaldi et al., 2024). Among them, the learned harmonic mean estimator (McEwen et al., 2021) has been shown to provide robust estimates of the Bayesian evidence in a variety of scenarios, both in likelihood- and simulation-based inference (McEwen et al., 2021; Polanska et al., 2023, 2024; Spurio Mancini et al., 2023). In particular, Polanska et al. (2024) integrated normalizing flows into the learned harmonic mean framework to learn the internal importance sampling target distribution, enhancing the robustness and scalability of the estimator. In contrast to the learned harmonic mean, most alternative approaches (Srinivasan et al., 2024; Rinaldi et al., 2024) compute the evidence for a surrogate of the posterior, limiting their accuracy. Gaussianized bridge sampling (GBS), introduced by Jia & Seljak (2020), is similar in spirit to the learned harmonic mean in that requires only posterior samples and considers an internal learned proposal distribution. In fact, bridge sampling in general can be viewed as a generalization of the standard importance sampling and original harmonic mean approaches to computing the evidence (Gronau et al., 2017). Nevertheless, restricting to the special case of the harmonic mean has numerous advantages. In addition to posterior samples, bridge sampling requires a second set of samples from the proposal distribution. Furthermore, the likelihood must be evaluated at these additional samples, which can add significant computational cost, and the proposal density must also be evaluated at all sample points, both those from the posterior and the proposal. Finally, use of the optimal bridge function (Meng & Wong, 1996) necessitates an iterative approach, again adding to the computational cost. The GBS approach suffers from all of these issues, whereas the learned harmonic mean does not. The primary disadvantage of restricting bridge sampling to the case of the harmonic mean is that the internal importance sampling target distribution must have thinner tails than the posterior, which is in any case already solved by the learned harmonic mean approach (McEwen et al., 2021). Another method to compute the Bayes factor is the Savage–Dickey density ratio (Dickey, 1971; Trotta, 2007), which is however limited to the case of nested models and requires the normalized marginal posterior to be computed.
@@ 3.2 Differentiable and probabilistic programming:
Modern ML ecosystems, such as JAX and PyTorch, are constructed with this in mind, requiring that each primitive operation within their libraries is coupled with associated gradient rules. In this way, using the chain rule one can construct more complex operations, which are automatically differentiable (Bartholomew-Biggs et al., 2000; Margossian, 2019). By automatic differentiation gradients can then be computed efficiently and accurately for complex operations, replacing numerical approximations of the derivatives, which are typically slow and unstable. Primarily due to its relative maturity, PyTorch provides a larger bank of primitive operations. However, JAX provides a more flexible ecosystem that is functional in nature (Wadler, 1992), avoiding boilerplate code that can arise in other ML-oriented frameworks, and being an efficient replacement for NumPy (Harris et al., 2020); it is thus well suited to the implementation of general physical models. In JAX complex operations may more easily be expressed, making it a natural ecosystem for the integration of differentiable programming in e.g. cosmological models (Campagne et al., 2023), cosmological simulations (Li et al., 2024) or underlying mathematical methods, such as spherical harmonic transforms (Price & McEwen, 2024).
Recently, probabilistic programming languages (PPLs) have been constructed leveraging these differentiable programming ecosystems. Through PPLs, statistical operations such as sampling and conditioning can be expressed as one-line statements, whilst also inheriting the differentiability of the underlying ecosystems upon which they are constructed. For instance, PPLs allows one to easily formulate Bayesian hierarchical models, which are widely used in cosmology (see e.g. Mandel et al., 2011; Shariff et al., 2016; Alsing et al., 2016, 2017; Hinton et al., 2019; Porqueres et al., 2021; Mandel et al., 2022; Porqueres et al., 2023; Loureiro et al., 2023; Sellentin et al., 2023; Kostić et al., 2023; Nguyen et al., 2024).
In this work we use NumPyro (Phan et al., 2019) [^11] https://num.pyro.ai/en/stable/, the NumPy backend for Pyro (Bingham et al., 2019). NumPyro is well-documented and easily interfaces with JAX, thus providing a mature framework to build a gradient-based inference pipeline in Python. We implement the likelihood for the problems considered (see Sect. 4 and Sect. 5) in the NumPyro PPL, which requires a separate likelihood implementation. Critically, the likelihood must be differentiable: CosmoPower-JAX provides emulation of the matter power spectrum that is differentiable, and we couple it with jax-cosmo (Campagne et al., 2023) [^12] https://github.com/DifferentiableUniverseInitiative/jax_cosmo for differentiable cosmological models to simulate observable spectra.
@@ 3.3 Scalable MCMC sampling: NUTS
HMC is a gradient-based sampling algorithm that is scalable to high-dimensional parameter spaces and does not suffer from the random-walk shortcomings of Metropolis–Hastings algorithms. NUTS further improves on the effective sample size relative to the gradient evaluation of HMC by preventing the sampling trajectory from returning to previously-visited regions of the parameter space. Though it requires a higher number of model evaluations, NUTS typically leads to more reliable convergence, and is thus ideal for sampling in the context of Stage IV surveys.
Traditionally, the gradients required for HMC sampling are computed by finite differences, which is computationally costly and inaccurate. Since we have a differentiable pipeline, gradients can be computed by automatic differentiation, providing acceleration and improved accuracy. In the future, alternative samplers (Gabrié et al., 2022; Wong et al., 2023; Karamanis et al., 2022a, b), or new samplers that further improve the sampling efficiency and thus lead to even more robust inference, can be straightforwardly incorporated into this paradigm.
@@ 3.4 Decoupled and scalable Bayesian model comparison: learned harmonic mean
Figure 1: Corner plot for the $37$-dimensional $\Lambda$CDM model, showing the posterior contours obtained with CosmoPower-JAX in red and the concentrated flow with temperature $T=0.8$ in blue. The $A_{\mathrm{s}}$ parameter indicates $\ln\left(10^{10}A_{\mathrm{s}}\right)$.
@@
4 Cosmic shear analysis
with 37 and 39 parameters
@@ 4.3 Computational approaches
Second, we compare this “traditional“ approach with our “future” paradigm, which replaces CAMB with the CosmoPower-JAX emulator, rewrites the likelihood in the auto-differentiable language JAX, making use of jax-cosmo, samples the posterior distribution with NUTS implemented in NumPyro, accelerated by automatic differentiation, and computes a robust estimate of the evidence from posterior samples using the learned harmonic mean implemented in harmonic. We run the NUTS chains in parallel on 12 A100 80GB GPUs, collecting 60 chains of 2000 samples each (after 400 samples of warm-up) for both $\Lambda$CDM and $w_{0}w_{a}$. As in Piras & Spurio Mancini (2023), parallelization is obtained by running multiple chains on each GPU. For harmonic, we train a rational-quadratic spline flow (Durkan et al., 2019), consisting of 4 layers and 64 spline bins, on a single GPU. We separate the NUTS samples into a training set the flow is trained on and an inference set that is used for estimation. We use $30\%$ of chains for training and the remaining $70\%$ for inference to compute the evidence on 48 CPUs with temperature $T=0.8$.
@@ 4.4 Results
The results are reported in Table 1(a). The values of the evidence when using CAMB $+$ nested sampling and CosmoPower-JAX $+$ NUTS $+$ harmonic use different implementations of the likelihood (the latter implemented in a probabilistic programming framework), including different normalizations, so their values cannot be compared directly. Nevertheless, the $\log$ BFs are in good agreement between the different approaches, with $0.78\pm 0.79$ for nested sampling and $1.53\pm 0.07$ for harmonic. Note that the evidence values correctly favour the ground truth model considered for the simulated data. While the evidence computed by the naïve flow estimator is also in agreement with a value of $1\pm 6$, notice that its error is considerably larger than the other estimators, as anticipated. While the evidence values computed by the “traditional” and “future” paradigms are in close agreement, computational times are dramatically different, with the “future” paradigm being two orders of magnitude faster, taking 2 days (with only 12 additional minutes required to compute the evidence) instead of roughly 8 months. Note that with respect to the analysis made by Piras & Spurio Mancini (2023), the computation times that we report here are higher since they include two models ($\Lambda$CDM and $w_{0}w_{a}$), as well as more chains to obtain a robust estimate of the evidence from harmonic. We also tested that running PolyChord in combination with the CosmoPower emulator yields consistent results, while taking about 16 days on 48 CPU cores (namely 2 runs of 8 days each to estimate the Bayes factor). We additionally report that a single evaluation of the log-likelihood gradient in forward (reverse) mode with our approach takes about 335 (80) seconds on a single CPU, and 15 (20) seconds on a single GPU, demonstrating the acceleration provided by modern hardware.
The contours for the posterior samples (red) alongside the concentrated flow at $T=0.8$ (blue) used for inference for this analysis are shown in Fig. 1 and Fig. 2 for the $\Lambda$CDM and $w_{0}w_{a}$ models respectively. A requirement for the learned harmonic mean estimator is that the concentrated flow is contained within the posterior. While it can be informative to inspect corner plots to ensure that, it is useful to consider several additional sanity checks, especially in high-dimensional contexts. In particular, we confirm that the values of the error estimate, kurtosis and the ratio between the square root variance of variance and variance estimates, introduced in detail by McEwen et al. (2021), are in agreement with the theoretical expectations.
@@ 5.2 Results
All numerical results for this higher dimensional analysis are reported in Table 1(b). The high number of parameters and the need to consider multiple probes pose a significant challenge for “traditional” methods, despite being quite realistic for next-generation surveys. We estimated that the computation of the Bayes factor in such a scenario would require about 12 years on 48 CPU cores, making it effectively impossible to perform model selection in a reasonable time. While an emulator could speed up nested sampling, it is unlikely that it could scale the analysis to such high-dimensional scenarios, given the relatively large variance of the “traditional” method already shown in Sect. 4. In our experiments, we found that even the 3x2pt analysis of a single survey with 57 parameters combining an emulator with PolyChord (with default hyperparameters) would struggle to converge in a reasonable time for the likelihood considered here.
On the other hand, the proposed “future” paradigm provides an estimate of the Bayes factor based on 40 (45) converged chains — with 1500 samples each, after 1000 samples of warm-up — for the $\Lambda$CDM ($w_{0}w_{a}$) model, totalling 8 days on 24 GPUs, with only 17 additional minutes required to compute the evidence. In this case, the flow model we consider is constructed from 2 layers and 128 spline bins. The $\log$ Bayes factor computed by harmonic is $1.9^{+0.7}_{-0.5}$, which correctly favours the ground truth model considered for the simulated data. The evidence computed by the naïve flow estimator is $2\pm 73$: its error is very large, as anticipated, rendering this approach unusable in practice. Again, the computation times reported here are higher than what presented in Piras & Spurio Mancini (2023) since we are now performing model comparison, and a higher number of chains was required to satisfy the diagnostics provided by harmonic.
The contours plots for the posterior samples (red) alongside the concentrated flow at $T=0.8$ (blue) used for inference for this analysis are shown in Fig. 3 and Fig. 4 for the $\Lambda$CDM and $w_{0}w_{a}$ models respectively. Fig. 5 and Fig. 6 show the one-dimensional marginal plots of the posterior samples (red) alongside the concentrated flow at $T=0.8$ (blue), for all the model parameters for $\Lambda$CDM and $w_{0}w_{a}$ respectively. As in Sect. 4.4, we ensure that we satisfy the sanity checks introduced by McEwen et al. (2021).